OpenStudy (anonymous):

write the equation of the line in slope-intercept form that passes through the given points (4,5) and (8,8)

OpenStudy (the_fizicx99):

Find the slope, M = y2 - y1 / x2 - x1 (x1,y1) and (x2,y2) M = 8 - 5 / 8 - 4 = 3/4 (I'd keep a fraction because it would make it easier to see it's rise 3 run 4) y - 5 = 3/4(x -4) y - 5 = 3/4x - 3 +5 +5 y = 3/4x + 2

OpenStudy (anonymous):

oh my gosh , its so confusing ;0 can you help me with just 1 more please..

OpenStudy (the_fizicx99):

Sure :>

OpenStudy (anonymous):

thank youuu.!

OpenStudy (the_fizicx99):

Cx

OpenStudy (anonymous):

(-2,5) and (3,-5)

OpenStudy (the_fizicx99):

Can you find the slope first? :O

OpenStudy (the_fizicx99):

*cough* \(\ \sf M = \dfrac{y2 - y1}{x2-x1} \) cx

OpenStudy (anonymous):

ohkay hold on ;o

OpenStudy (the_fizicx99):

 Artist ~